OpenStudy (anonymous):

Lucia and Randy have signed a contract to purchase a home. The closing date is March 27, and the buyer owns the property on the day of closing. The selling p rice of the home is $852,500. Lucia and Randy obtained a fixed-rate mortgage from a bank for $665,000 at 7.55% interest. The seller has already paid $15,378.15 in property taxes for the coming year. How much will Lucia and Randy owe in prorated expenses?

OpenStudy (anonymous):

@phi can you help me with this one?

OpenStudy (phi):

I am not an accountant, but I assume they will have to repay the owner for some of the taxes. March 26 is the 85th day of the year (for a non-leap year) so the original owner owes 85/365 of the taxes and L & R owe (365-85)/365 of the taxes.

OpenStudy (phi):

They will also owe $852,500 - $665,000

OpenStudy (anonymous):

so 76.7% of the taxes plus 187500?

OpenStudy (phi):

owe in prorated expenses I assume that is just the taxes. The difference between the price and loan is not a pro-rated expense (is it ?)

OpenStudy (anonymous):

no?

OpenStudy (phi):

I don't study this stuff, just using common sense. Pro-rated means to divide, assess, or distribute (something) proportionately They do that with the taxes, but not with the difference between the price and loan

OpenStudy (anonymous):

right. That makes sense.

OpenStudy (anonymous):

so the answer should be 11795.04

OpenStudy (phi):

I don't know how many days they use for a full year. It might be 365.25 or it might be 365 or 366 depending on if its a leap year. Or they might always use 365. It will make a small difference.

OpenStudy (anonymous):

I think that is right. I have another question it is different. Do you think you could help me?

OpenStudy (phi):

what is it ?

OpenStudy (anonymous):

4. Jerry is purchasing a condominium for $325,500 and has financed 90% of it. He has purchased 3 discount points, has to pay a 1.65% origination fee and a brokerage fee of $100 plus 0.25% of the loan. What are his total fees that he must pay? (1 point) $15,461.25 $16,293.64 $14,454.56 $15,317.13

OpenStudy (anonymous):

I thought the answer was 16,293 but the answer is 14454.56 but I am not sure how they got that or what I did wrong.

OpenStudy (phi):

loan amount is 90% of 325,500

OpenStudy (anonymous):

yes,

OpenStudy (phi):

loan amt is 292,950 3 discount points Each discount point generally costs 1% of the total loan I get 0.03*292,950= 8,788.50 has to pay a 1.65% origination fee 0.0165*Loan= 4833.66 brokerage fee of $100 0.25% of the loan= 732.38

OpenStudy (anonymous):

I did 325500*.9=292950 then I did 292950*.0025=735.375 +100= 832.375. Then I did 1.65+3=4.65 then 325500*.0465=15135.75 Then I added and got 1535.75+832.38=15968.13

OpenStudy (anonymous):

Oh. The discount points are from the cost of the loan not the total cost!

OpenStudy (anonymous):

Thank you. I was doing everything on the original cost not the cost of the loan! Thank you for your help!

OpenStudy (phi):

if you were just borrowing 10% of the cost (you saved the rest and don't need a big loan) you surely would not want to pay fees on the original price, just the loan.